Seminar Example

Seminar Example

T. W. Shield

Department of Aerospace Engineering and Mechanics University of Minnesota November 22, 2002

Seminar Example

All LaTeX stuff as usual

• itemize

• enumerate

• etc...

Seminar Example

Graphics also work as usual

\epsfslidesize

\includegraphics[scale=0.45]{figure.eps}

Seminar Example

Minipage Environment

• Useful for Two Columns of • Bullets

• etc.

• Red • Blue • Green